OpenStudy (anonymous):

Take the limit of this trig function:

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi/4}\frac{ 7-7tanx }{ sinx-cosx }\]

OpenStudy (anonymous):

use L'Hopital's rule: \[\frac{-7\sec^2(x)}{\cos(x)+\sin(x)}\]

OpenStudy (anonymous):

take the limit as x goes to pi/4: \[\frac{-7\sec^2(\pi/4)}{\cos(\pi/4)+\sin(\pi/4)} = \frac{-7(2)}{\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}}}\]

OpenStudy (anonymous):

-7/sqrt(2)?

OpenStudy (anonymous):

No. That can't be right. I must have made a mistake in my algebra.

OpenStudy (anonymous):

I'm lost. -7/sqrt(2) is not the right answer.

OpenStudy (anonymous):

I didn't use l'hopital's rule because we haven't gotten there yet.

OpenStudy (anonymous):

I multiplied top and bottom by cosx

OpenStudy (anonymous):

and pulled out the 7

OpenStudy (anonymous):

so I was left with \[\frac{ cosx-sinx }{ sinx-cosx }\]

OpenStudy (anonymous):

sorry, that's wrong

OpenStudy (anonymous):

should be a cosx multiplied in the denominator

OpenStudy (anonymous):

\[\frac{ cosx-sinx }{ cosx(sinx-cosx) }\]

OpenStudy (anonymous):

\[\frac{2}{\frac{1}{\sqrt{2}}+\frac{1}{\sqrt{2}}} = \sqrt{2}\]

OpenStudy (anonymous):

oh, so it's just \[7\sqrt{2}\]

OpenStudy (anonymous):

the answer is: \[7\sqrt{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Hmm, it's telling me that the answer is incorrect.

zepdrix (zepdrix):

 Hmm this is what i'm getting